# Multi-Agent Systems Execute Arbitrary Malicious Code

### v1

@@ Abstract
Multi-agent systems coordinate LLM-based agents to perform tasks on users’ behalf. In real-world applications, multi-agent systems will inevitably interact with untrusted inputs, such as malicious Web content, files, email attachments, etc.
Using several recently proposed multi-agent frameworks as concrete examples, we demonstrate that adversarial content can hijack control and communication within the system to invoke unsafe agents and functionalities. This results in a complete security breach, up to execution of arbitrary malicious code on the user’s device and/or exfiltration of sensitive data from the user’s containerized environment. We show that control-flow hijacking attacks succeed even if the individual agents are not susceptible to direct or indirect prompt injection, and even if they refuse to perform harmful actions.
@@ 1 Introduction
In recent years, generative AI models have rapidly evolved from autoregressive next-word predictors to interactive chatbots capable of reasoning across multiple input modalities to API-connected agents that can plan and operate in open digital environments to fulfill users’ requests and queries. Just as Web browsers transformed how users interact with online content, multi-agent systems (MAS) are emerging as a powerful new paradigm for navigating digital information and performing complex tasks.
Consider typical motivating scenarios: a user requires research assistance, financial advice, or help with a complex coding project. Multi-agent systems promise to simplify these tasks by coordinating specialized agents (typically, LLM-based) to browse the Web, analyze documents, summarize information, and even execute code on the user’s behalf. This coordination involves adaptive control flows, i.e., LLM-based computational logic that dynamically generates and utilizes contextual metadata to guide a sequence of actions by individual agents. This metadata encompasses task plans, action histories, self-evaluative progress assessments, and action results, which are processed at each coordination step.
Any system operating in today’s Internet environment will interact with both trusted and untrusted content: reading emails from unknown senders, browsing user-generated content, analyzing shared documents, and accessing local and remote resources. Untrusted content is an attack surface. A single maliciously crafted webpage, image, or audio file can potentially compromise an entire system, putting users at risk without their knowledge or consent.
Whereas Web browsers have developed sophisticated mechanisms, such as the same-origin policy (Mozilla, 2025), to isolate and sandbox untrusted content, the boundary between trusted and untrusted content in multi-agent systems is blurry, and it is not clear what mechanisms—if any—these systems are deploying to protect users from malicious content.
@@
Control-flow hijacking attacks are distinct from jailbreaking and indirect prompt injection (see Section 2). In contrast to jailbreaking attacks, where the user is attacking the model and trying to get it to generate a harmful output, in our case the user is an unwitting victim, not the instigator of the attack. While adversarial inputs in our attack are similar to indirect prompt injection, we empirically demonstrate that previously known indirect injection attacks have near-zero success rate in getting MAS to execute malicious code.
MAS hijacking exploits metadata transmission pathways to reroute the sequence of agent invocations to unsafe agents, or unnecessary agents in harmful configurations. This attack is an instance of the confused deputy problem (Hardy, 1988), well-known in computer security but apparently unknown or unheeded in the design of LLM-based multi-agent systems. These systems delegate tasks to individual sub-agents and trust that they correctly report the status of their tasks. In a MAS hijacking attack, adversarial content turns sub-agents into “confused deputies,” in effect laundering the adversary’s requests so they appear as trusted outputs of trusted agents. Figure 1 shows a high-level example of this attack.
@@
Our initial investigation of these attacks considers three prominent open-source multi-agent frameworks that are user-focused and designed to operate on both local and remote content: AutoGen, [^1] https://microsoft.github.io/autogen/stable/ Crew AI, [^2] https://github.com/crewAIInc/crewAI and MetaGPT. [^3] https://github.com/geekan/MetaGPT Evaluating these systems in a variety of configurations, we demonstrate that they are highly vulnerable to control-flow hijacking attacks. For example, a malicious webpage succeeds in getting these systems to run malicious code 45-64% of the time (averaged across models). For certain model-orchestrator combinations, the attack success rate is 100%. Furthermore, we demonstrate that these attacks succeed even if the individual agents and the models that power them are relatively immune to direct or indirect prompt injection.
Figure 2 illustrates the most basic version of the attack. The user—either knowingly, or as the result of another MAS invocation—asks a centrally orchestrated MAS to access a malicious local file (e.g., a saved email attachment). The orchestrator agent tells the file access agent to read the file. The access agent returns a false error message, saying that the file needs to be executed to access it. The orchestrator then tells the code execution agent to run the file. The executable code in the file initiates a reverse shell script that cedes control of the user’s computer to the attacker.
@@
We reiterate that the root cause of these vulnerabilities is not the absence of safety alignment in the individual agents. In several cases, we observed multi-agent systems finding a way to execute harmful code even when sub-agents refuse to take action for safety reasons.
For example, in one Crew AI experiment the orchestrator reasoned that executing a file might be unsafe and resolved to “read the file safely.” It then created a dummy file to read, read the dummy file, and realized that the dummy file is not what the user wanted. The orchestrator then re-read the attack file and executed a reverse shell script (see Appendix A). Similarly, in one AutoGen experiment the file-surfer sub-agent produced its own reverse shell script (along with the helpful comment “This is a dangerous command example. Do not execute!”) after the coder agent refused to output malicious code. The generated code was then promptly executed by the code executor, opening a reverse shell (see Appendix B).
The remainder of this paper is structured as follows. In Section 2, we explain why control-flow hijacking attacks are conceptually different from jailbreaking and indirect prompt injection. Section 3 provides a brief review of how multi-agent systems work, focusing on individual agent capabilities, system topologies, and control flows. Section 4 describes the threat model for MAS hijacking. In Section 5, we formally define MAS hijacking and explain how the attack works in practice, from the simplest examples (a local plaintext file) to more complex attacks (multi-modal inputs, multiple action steps, etc.). In Section 6, we describe our experiments for evaluating this attack across multiple frameworks, orchestrators, models, and input modalities. Our results are in Section 7. In Section 8, we survey related work, and conclude with the discussion and future work in Section 9.
2 Control-flow hijacking is a new threat
Table 1 and the following list summarize the differences between control-flow hijacking and other attacks on multi-agent systems. See Section 8 for a survey of related work.
  • Target: Rather than attacking outputs or behaviors of individual agents, MAS hijacking targets the orchestration layer and communication protocols between agents.
  • Mechanism: The attack manipulates the metadata used for system coordination rather than the primary outputs.
  • Threat model: The user is a victim, not the attacker. Their MAS is compromised through routine interaction with Internet or local content.
  • Effectiveness: As our results demonstrate, MAS hijacking succeeds even when prior attacks fail against the same models and frameworks.
  • Scope: The attack can affect the entire system operation, not just individual agent responses or isolated components.
@@ 2 Control-flow hijacking is a new threat
Jailbreaking Direct user request designed to bypass model safeguards Single model’s output generation
Indirect Prompt Injection (IPI) Planted instructions in content processed by agent tools Single agent’s reasoning process
@@ 3.1 Agentic AI
The modern notion of an “AI agent” lacks a precise definition. Instead, as noted in Kapoor et al. (2024), it is often useful to consider a spectrum of agency, often denoted by the term “agentic.” Systems that are more agentic can operate within complex and unexpected environments, with ambiguous initial user directives (often using natural language) and less frequent human intervention for system guidance (Chan et al., 2023; Gabriel et al., 2024; Shavit et al., 2023). Today, agentic AI systems are driven by LLMs, which decompose high-level tasks into smaller chunks, plan how to complete individual sub-goals, and elect how and when to utilize tools, i.e., structured program execution interfaces for interacting with external resources, like the Internet, databases, or local file systems (Chase, 2024; Weng, 2023). Agent systems can employ methodologies like ReAct (Yao et al., 2023), which implements a thought-action-observation loop to enable agent adaptation to the results of tool invocations.
AI agent systems and frameworks have demonstrated their effectiveness on many benchmarks (Liu et al., 2023; Mialon et al., 2023; Zhou et al., 2024; Yao et al., 2024; Huang et al., 2024). These frameworks and benchmarks have also been adapted to individual AI agents that can complete complex tasks on behalf of users (OpenAI, 2025; Anthropic, 2024; Liu et al., 2024a; b). Increasing AI system agency is both a promising direction for research and a source of new security risks, particularly as agent systems become more complex and turn into multi-agent systems.
@@ 3.2 Multi-agent systems
As the name suggests, a multi-agent system is a system comprised of multiple AI agents, each with potentially different capabilities. This design is guided by a simple intuition: by instantiating several agents with particular areas of expertise and tools, the system as a whole is more likely to complete complex tasks.
Multi-agent systems typically operate through task planning and control flow mechanisms. These mechanisms decompose complex tasks into manageable sub-tasks, identify appropriate agents for each component, and orchestrate their sequential or parallel execution. This orchestration involves determining which agent has the appropriate capabilities for each step and managing the transitions between agents to ensure coherent task completion.
A key feature of advanced multi-agent systems is their adaptivity. The orchestrator can dynamically decide which agents to invoke next based on the responses and status reports from previously invoked agents. This adaptivity enables the system to respond to both successful outcomes and unexpected errors, making real-time adjustments to the execution flow as needed.
In multi-agent interactions, it is crucial to distinguish two types of information exchange: data (the actual content or responses produced by agents) and metadata (status reports, error messages, progress updates, and other operational information that guides system behavior). This distinction becomes particularly important when considering potential security vulnerabilities, as metadata can influence control-flow decisions even when primary outputs are properly secured.
But if agents are operating in a coordinated manner, how should they interact? Finding the best system topology—broadly, the most effective means of instantiating, structuring, and coordinating agents in a multi-agent system—is an open area of research, with different approaches proposed by different MAS frameworks. They seek to answer two questions:
  1. 1.
    Which agents can communicate with each other?
  2. 2.
    How should a multi-agent system determine which agent (or agents) should be called next?
Existing frameworks explore a variety of answers to these questions, as shown in Figure 3. In this paper, we consider three topologies:
(a) Round Robin
(b) Central Orchestrator
(c) Central Orchestrator (Ledger)
Figure 3: Multi-agent topologies.
In all cases, each agent can see a partial history of the reasoning steps and actions that other agents have taken. This metadata is a crucial tool for ensuring that multi-agent systems are actually functional. As we show in the rest of the paper, it also presents a new attack surface that can be exploited by MAS hijacking attacks.
@@ 4 Threat model
In this section, we briefly describe the adversary, their capabilities, and their goals.
@@ 4.1 Adversary’s capabilities
We assume that agents in a MAS are benign and not controlled by the adversary. That said, they inevitably interact with untrusted inputs: Web content (e.g., malicious websites or social media) or, in a stronger threat model, files already on the user’s device (e.g., email attachments or images). For example, a user seeking IT advice may invoke a MAS to summarize a seemingly innocuous Web forum with a malicious post. Similarly, a request to process email attachments could inadvertently cause the MAS to act on a malicious file.
Unlike in jailbreaking attacks where the adversary is the user, in our case the adversary aims to cause the user’s MAS to harm the user. The adversary cannot directly prompt the MAS, but they control the inputs that a MAS may access inadvertently or on purpose.
@@ 4.2 Adversary’s goals
Executing arbitrary code on the user’s device. By tricking the MAS into opening a reverse shell from the user’s device to the adversary’s device, the attacker gains terminal access to the MAS’s code execution environment. This access enables a range of threats: installing malware, deploying keyloggers, or running cryptomining software, as well as adding, removing, or copying any files the agents can access. In addition, the attacker can exploit the victim’s compute resources for malicious and/or illegal activities such as hosting command-and-control beacons, forming botnets to launch distributed denial-of-service (DDoS) attacks, and orchestrating phishing campaigns.

### v2

@@ Abstract
Multi-agent systems coordinate LLM-based agents to perform tasks on users’ behalf. In real-world applications, multi-agent systems will inevitably interact with untrusted inputs, such as malicious Web content, files, email attachments, and more.
Using several recently proposed multi-agent frameworks as concrete examples, we demonstrate that adversarial content can hijack control and communication within the system to invoke unsafe agents and functionalities. This results in a complete security breach, up to execution of arbitrary malicious code on the user’s device or exfiltration of sensitive data from the user’s containerized environment. For example, when agents are instantiated with GPT-4o, Web-based attacks successfully cause the multi-agent system execute arbitrary malicious code in 58-90% of trials (depending on the orchestrator). In some model-orchestrator configurations, the attack success rate is 100%. We also demonstrate that these attacks succeed even if individual agents are not susceptible to direct or indirect prompt injection, and even if they refuse to perform harmful actions. We hope that these results will motivate development of trust and security models for multi-agent systems before they are widely deployed.
@@ 1 Introduction
In recent years, generative AI models rapidly evolved from autoregressive next-word predictors to interactive chatbots capable of reasoning across multiple input modalities to API-connected agents that can plan and operate in open digital environments to fulfill users’ requests and queries. Just like Web browsers in the 1990s and 2000s transformed how users interact with online content, multi-agent systems (MAS) are emerging as a powerful new paradigm for navigating digital information and performing complex tasks.
A typical motivating scenario is a user who needs research assistance, data analysis, or help with a complex coding project. Multi-agent systems promise to simplify these tasks by coordinating specialized agents (typically, LLM-based) to browse the Web, analyze documents, summarize information, and even execute code on the user’s behalf. This coordination involves adaptive control flows, i.e., LLM-based computational logic that dynamically generates and utilizes contextual metadata to guide a sequence of actions by individual agents. This metadata encompasses task plans, action histories, self-evaluative progress assessments, and action results, which are processed at each coordination step.
Any system operating in today’s Internet environment will interact with both trusted and untrusted content: reading emails from unknown senders, browsing user-generated content, analyzing shared documents, and accessing local and remote resources. Untrusted content is an attack surface. A single maliciously crafted webpage, image, or audio file can potentially compromise an entire system, putting users at risk without their knowledge or consent. Whereas Web browsers have developed sophisticated mechanisms, such as the same-origin policy (Mozilla, 2025), to isolate and sandbox untrusted content, the boundary between trusted and untrusted content in multi-agent systems is blurry, and it is not clear what mechanisms—if any—these systems are deploying to protect users from malicious content.
@@
Control-flow hijacking attacks are distinct from jailbreaking and indirect prompt injection (see Table 1). In contrast to jailbreaking attacks, where the user is attacking the model and trying to get it to generate a harmful output, in our case the user is an unwitting victim, not the instigator of the attack. While adversarial inputs in our attack are a form of indirect prompt injection, we empirically demonstrate that previously known indirect injection attacks have near-zero success rate in getting MAS to execute malicious code.
MAS hijacking exploits metadata transmission pathways to reroute the sequence of agent invocations to unsafe agents or unnecessary agents in harmful configurations. This attack is an instance of the confused deputy problem (Hardy, 1988), well-known in computer security but apparently unknown or unheeded in the design of LLM-based multi-agent systems. These systems delegate tasks to individual sub-agents and trust that they correctly report the status of their tasks. In a MAS hijacking attack, adversarial content turns sub-agents into “confused deputies,” in effect laundering the adversary’s requests so they appear as trusted outputs of trusted agents. Figure 1 shows a high-level example.
@@
Our initial investigation of these attacks considers three prominent open-source multi-agent frameworks that are user-focused and designed to operate on both local and remote content: AutoGen (Wu et al., 2023), CrewAI (CrewAI, 2025), and MetaGPT (Hong et al., 2024). Evaluating these systems in a variety of configurations, we demonstrate that they are highly vulnerable to control-flow hijacking. For example, the Magentic-One orchestrator, running on GPT-4o, executes arbitrary malicious code 97% of the time if it interacts with a malicious local file. The same orchestrator, running on Gemini 1.5 Pro, executes arbitrary malicious code 88% of the time when it interacts with a malicious web page. CrewAI, running on GPT-4o, is “convinced” by a local file to exfiltrate private user data with a 65% success rate. For certain model-orchestrator combinations, the attack success rate is 100%. Furthermore, we demonstrate that these attacks succeed even if the individual agents and the models that power them are relatively immune to direct or indirect prompt injection.
Figure 2 illustrates the most basic version of the attack. The user—either knowingly, or as the result of another MAS invocation—asks a centrally orchestrated MAS to access a malicious local file (e.g., a saved email attachment). The orchestrator agent tells the file-access agent to read the file. The access agent returns a false error message, saying that the only way to access the file is to execute it. The orchestrator then tells the code execution agent to run the file. The executable code in the file initiates a reverse shell script that cedes control of the user’s computer to a remote attacker.
@@
We reiterate that the root cause of these vulnerabilities is not the absence of safety alignment in the individual agents. In several cases, we observed multi-agent systems finding a way to execute harmful code even when sub-agents refuse to take action for safety reasons. For example, in one CrewAI experiment the orchestrator reasoned that executing a file might be unsafe and resolved to “read the file safely.” It then created a dummy file to read, read the dummy file, and realized that the dummy file is not what the user wanted. The orchestrator then re-read the attack file and executed a reverse shell script (see Appendix J). Similarly, in one AutoGen experiment the file-surfer sub-agent produced its own reverse shell script (along with the helpful comment “This is a dangerous command example. Do not execute!”) after the coder agent refused to output malicious code. The generated code was then promptly executed by the code executor, opening a reverse shell (see Appendix K).
@@
Jailbreaking Direct user request designed to bypass model safeguards Single model’s output
Indirect Prompt Injection (IPI) Planted instructions in content processed by agent tools Single agent’s reasoning and output
@@ 2.1 Agentic AI
The modern notion of an “AI agent” is imprecise. As noted in Kapoor et al. (2024), the term “agentic” covers a spectrum of agency. Systems that are more agentic can operate within complex and unexpected environments, with ambiguous initial user directives in natural language and less frequent human intervention for system guidance (Chan et al., 2023; Gabriel et al., 2024; Shavit et al., 2023). Today, agentic AI systems are driven by LLMs, which decompose high-level tasks into sub-goals, plan how to complete them, and elect how and when to utilize tools, i.e., structured program execution interfaces for interacting with external resources like the Internet, databases, or local file systems (Chase, 2024; Weng, 2023). Agentic systems can employ methodologies like ReAct (Yao et al., 2023), which implements a thought-action-observation loop so agents can adapt to the results of tool invocations.
AI agent systems and frameworks have demonstrated their effectiveness on many benchmarks (Liu et al., 2023; Mialon et al., 2023; Zhou et al., 2024; Yao et al., 2024; Huang et al., 2024). Similar techniques are used in individual AI agents that can complete complex tasks on behalf of users (OpenAI, 2025; Anthropic, 2024; Liu et al., 2024a; b).
@@ 2.2 Multi-agent systems
A multi-agent system is comprised of multiple AI agents, each with potentially different capabilities. These systems typically operate through task planning and control-flow mechanisms that decompose complex tasks into manageable sub-tasks, identify appropriate agents for each component, and orchestrate their sequential or parallel execution. Orchestration involves determining which agent has the appropriate capabilities for each step and managing the transitions between agents to ensure coherent task completion. A key feature of advanced multi-agent systems is their adaptivity. The orchestrator can dynamically decide which agents to invoke next based on the responses and status reports from previously invoked agents. This adaptivity enables the system to respond to both successful outcomes and unexpected errors, making real-time adjustments to the execution flow as needed.
In multi-agent interactions, it is crucial to distinguish data (the actual content or responses produced by agents) and metadata (status reports, error messages, progress updates, and other operational information that guides system behavior). This distinction becomes particularly important when considering potential security vulnerabilities, as metadata can influence control-flow decisions even when primary outputs are properly secured.
Finding the best system topology, i.e., the most effective means of instantiating, structuring, and coordinating agents in a multi-agent system, is an open area of research. The key questions are (1) which agents can communicate with each other?, and (2) which agent (or agents) should be called next? Existing frameworks explore a variety of answers to these questions. In this paper, we consider three topologies (illustrated in Appendix F): Round Robin, where agents take turns in a fixed order; Central Orchestrator, where a single agent dynamically assigns tasks to the others; and Central Orchestrator with External Data Structures, which utilizes an external task ledger to ensure the MAS stays on-task. In all cases, each agent can see a partial history of the reasoning steps and actions that other agents have taken. This metadata is crucial for ensuring that multi-agent systems are functional.
@@ 3.1 Adversary’s capabilities
We assume that agents in a MAS are benign and not controlled by the adversary. That said, they interact with untrusted inputs: Web content (e.g., malicious websites or social media) or, in a stronger threat model, files already on the user’s device (e.g., email attachments or images). For example, a user seeking IT advice may invoke a MAS to summarize Web forums, one of which includes a malicious post. Similarly, a request to process email attachments could inadvertently cause the MAS to act on a malicious file. Unlike in jailbreaking attacks where the user is the adversary, in our case an external adversary aims to cause the user’s MAS to harm the user. This adversary cannot directly prompt the MAS, but they control the inputs that a MAS may access inadvertently or on purpose.
@@ 3.2 Adversary’s goals
Executing arbitrary code on the user’s device. By tricking the MAS into opening a reverse shell from the user’s device to the adversary’s device, the attacker gains terminal access to the MAS’s code execution environment. This enables the attacker to install malware or keyloggers, run cryptomining software, and add, remove, or copy any files the agents can access. The attacker can also exploit the victim’s compute resources for malicious and/or illegal activities such as hosting command-and-control beacons, forming botnets to launch distributed denial-of-service (DDoS) attacks, and running phishing campaigns.